# Top-Theta Attention: Sparsifying Transformers by Compensated Thresholding

### v1

@@ 5 Related Work
In the line of work on the content-based sparsity, the most fundamental approach is Top-k attention (Gupta et al., 2021), which selects the $k$ highest values in each row of the $A$ matrix. However, in practical kernel implementations, where the attention rows are tiled, it is highly difficult to implement the top-k search algorithm due to the full row dependency. Works such as Energon (Zhou et al., 2022) went further and aimed at reducing the reads from the K-cached matrix by iteratively quantizing K, computing the $A$ matrix, and keeping the higher values. However, reducing access to K prohibits Energon from accurately computing the $S$ matrix, which is therefore as limited as not using our compensation methods in Top-{greek}j. The SpAtten (Wang et al., 2021) and A3 (Ham et al., 2020) present Top-k attention variants which require specialized hardware for accelerating the top-k search, which can be avoided by Top-{greek}j. The SparQ (Ribar et al., 2023) work also relies on the top-k search, however, it performs a similar compensation to our VMC to compensate for the rows of V matrix, which did not participate in the SV product ( 4) due to the sparsification of the attention. The Swin Transformer (Liu et al., 2021) focuses on a vision transformer (i.e., no autoregressive decoding) applying a fixed sparsity pattern, where each token (patch) attends to its neighboring M×M patches.
Learned Token Pruning (LTP) (Kim et al., 2022) is closely related to our work as it discards tokens in the attention matrix based on a model-trained threshold. While LTP benefits from the simplicity of the thresholding approach, it requires full model retraining. Our method shares the threshold concept but improves upon it by identifying an inductive threshold within an already-trained model. This makes our approach more appealing for large models, where training or fine-tuning can be prohibitively expensive. Similarly, the work of Lou et al. (Lou et al., 2024) also targets the selection of top-k elements but requires retraining.
@@
Numerous works focused on fixed attention sparsity. For example, in Longformer (Beltagy et al., 2020), only a few manually picked “global attention” tokens are allowed to attend to other tokens, whereas other tokens are restricted to the neighboring tokens in a dilated sliding window. Our work, on the other hand, capitalizes on identifying content-based sparsity in the attention matrix, as it offers higher accuracy gains. The work of Kim and Cho (Kim & Cho, 2021) requires retraining, although accommodates dynamic shortening of the sequence on demand at inference time.
@@ 6 Conclusion
We have presented Top-{greek}j, a new sparse attention algorithm based on fixed and calibrated thresholds. At inference time, it boils down to a simple elementwise operation, which overcomes the limitation of full-row dependent algorithms, such as Top-k, and thereby unconstraining tiling and distributed inference implementations. Moreover, only a small subset of calibration samples is required to calibrate the thresholds, which we showed to be resilient to distribution shifts. Therefore, a short calibration is needed once per every model. Furthermore, we show minor to no performance degradation using 10× less attention elements at the computationally-bound prefill phase and using 3× less V matrix rows at the memory bandwidth-bound generative decoding phase. These reductions unlock promising speedups for inference.
@@ Appendix A Impact Statement
This paper presents work whose goal is to advance the field of efficient Machine Learning. All potential societal consequences are unrelated to the specific work but are related to machine learning applications in general.
Appendix B Threshold Calibration
In this appendix section, we show more calibrated threshold values (as a function of the sequence length) in more layers and heads.  Figure 12 visualizes the thresholds as a function of a sequence length that was calibrated for The LLaMA2-7b model for the Hellaswag dataset. Different heads have evidently different threshold values, hence the importance of per-head individual calibrations
@@ Appendix B Threshold Calibration
Appendix C Multi-k Cumulative Calibration
Algorithm 1, which we presented in the paper, calibrates thresholds of a single attention head for a single target $k$. We propose to generalize it to calibrate all thresholds for a broader range of target $k$ and to do it in one pass over the calibration samples. We call such a calibration procedure “Multi-k Cumulative” (MKC).
@@ Appendix C Multi-k Cumulative Calibration
Algorithm 2 describes how a threshold function $\theta(k)$ can be calibrated. First, for every given calibration sample, let $v$ represent the $r^{th}$ attention row. The $v$ undergoes a sorting and then is treated as a sequence of half-open intervals. For each interval, we treat its smaller endpoint as its threshold, and we also associate with it an effective $k$ (effective $k$ of a threshold w.r.t the vector $v$ is the number of elements that are greater than the threshold). For example, on line 2 we represent by $\langle r-1,[v_{1},v_{2})\rangle$ the interval between $v_{1}$ (inclusive) and $v_{2}$ (exclusive) that corresponds to the threshold $v_{1}$ and an effective $k$ of $r-1$. Second, per-calibration-sample interval sequences collected in the set $\Theta_{r}$ are merged to represent the function $\bar{k}(\theta):\mathbb{R}\rightarrow\mathbb{R}$ which maps each interval to a threshold and to an average effective $k$ achievable by this threshold on the calibration set. This merging is done by the MergeIntervals($\Theta_{r}$) subroutine, which computes the following:
@@
where we define $\Theta_{r,\theta}=\{t|\theta\in t_{1}\wedge t\in seq\wedge seq\in\Theta_{r}\}$ as a subset of intervals that include the $\theta$ in theinterval. Note that the resulting average effective $k$ might be fractional due to the averaging. Finally, the desired $\theta(k)$ is given by the inverse of $\bar{k}(\theta)$, at points where the function $\bar{k}(\theta)$ maps to a natural value. The full implementation of the MergeIntervals routine is available in the supplied source code, and we show its visualization on Figure 13.
@@
0:  $\mathcal{C}$ (Calibration set of inputs)
1:  $\Theta_{r}=\emptyset,\forall r$empty sets of observed interval sequences}
@@
3:     if is_prefill($X$then
4:        $A=Q(X)K^{T}(X)$
5:        $n=$NumRows$(A)$
6:        for $r=1$ to $n-1$ do
7:           $\bm{v}=$Sort$(A_{r})$
@@
Figure 14 presents how do the merged intervals ($\bar{k}(\theta)$) look like for a few selected layers and heads on LLaMA2-7b model, calibrated using MKC on the Hellaswag.
@@
Figure 14: MKC calibration of LLaMA2-7b on Hellaswag dataset. Right: the $\bar{k}_{r}(\theta)$ function as obtained from Algorithm 2 by merging the intervals for sequence length $r=800$. Middle: zoom into a range of interest of $\bar{k}_{r}^{-1}(k)$ around $k=64$, the selected threshold is marked on using $\theta_{r}(64)=\bar{k}_{r}^{-1}(64)$, Left: thresholds obtained from MKC for all sequence lengths (also other than $r=800$.
It is worth mentioning that the main downside of the multi-k calibration is that it is very time and memory-consuming to perform all these interval mergings individually for every layer, head, and attention row id (i.e., sequence length). Speeding up the MergeIntervals routine, using reasonable approximations (such as subsampling) is an interesting research direction. The second slight downside is that it does not allow to apply topk at calibration, as we used to apply in Algorithm 1, since during MKC there is a multitude of potential $k$ parameters we are targeting.
Appendix D V-Mean Compensation
In this section, we provide a formal proof for our V-Mean Compensation (VMC) serving as a good approximation of the sparsified $\tilde{\bm{s}}\tilde{V}$ product to the full $\mathbf{s}V$ product.
Lemma D.1.
Let $\bm{s}\in[0,1]^{n}$ denote the post-softmax attention vector, and let $\tilde{\bm{s}}\in[0,1]^{\tilde{k}}$ denote its thresholded variant, containing the $\tilde{k}$ selected values of $\bm{s}$, and let $\mathcal{I}\in\left\{0,\dots,n-1\right\}^{\tilde{k}}$ denote the indices of $\bm{s}$ that surpassed the threshold such that $\forall i\in\{0,\dots,\tilde{k}-1\}:\tilde{\bm{s}}_{i}=\bm{s}_{\mathcal{I}_{i}}$. Let $V\in\mathbb{R}^{n\times d}$ be the value matrix and let $\tilde{V}\in\mathbb{R}^{\tilde{k}\times d}$ be consisting only of the selected V rows such that $\forall i\in\{0,\dots,\tilde{k}-1\}:\tilde{V}_{i}=V_{\mathcal{I}_{i}}$. We claim that in the expectation the full product $\bm{p}=\bm{s}V\in\mathbb{R}^{d}$ is equal to the thresholded attention plus the residual probability mass $\beta$ (10) multiplied by the mean-V row $\bm{\mu}$ (9). Namely,
$\forall 0\leq j<d:E[\bm{p}_{j}]=(\bm{\tilde{s}}\tilde{V})_{j}+\beta\bm{\mu}_{j}$ (13)
@@ Appendix D V-Mean Compensation
$\begin{split}E[\bm{p}_{j}]&=E\Big{[}\sum_{i=0}^{n-1}s_{i}V_{ij}\Big{]}\\ &=E\Big{[}\sum_{i\in\mathcal{I}}s_{i}V_{ij}\Big{]}+E\Big{[}\sum_{i\in\bar{\mathcal{I}}}s_{i}V_{ij}\Big{]}\\ &=(\bm{\tilde{s}}\tilde{V})_{j}+\sum_{i\in\bar{\mathcal{I}}}E[s_{i}V_{ij}]\\ &\underset{s\perp\!\!\!\!\perp V_{j}}{=}(\bm{\tilde{s}}\tilde{V})_{j}+\sum_{i\in\bar{\mathcal{I}}}E[s_{i}]E[V_{ij}]\\ &\underset{s,V_{j}\text{uniform}}{=}(\bm{\tilde{s}}\tilde{V})_{j}+\sum_{i\in\bar{\mathcal{I}}}\frac{\beta}{n-\tilde{k}}\bm{\mu}_{j}\\ &=(\bm{\tilde{s}}\tilde{V})_{j}+(n-\tilde{k})\frac{\beta}{n-\tilde{k}}\bm{\mu}_{j}\\ &=(\bm{\tilde{s}}\tilde{V})_{j}+\beta\bm{\mu}_{j}\end{split}$ (14)
@@
    $s\perp\!\!\!\!\perp V_{j}$, that is the attention vector $\bm{s}$ is statistically independent on the elements in the columns of matrix $V$. They are conditionally independent given the input $X$ from which they were originally computed (1).
@@
    The distribution of $s_{i},\forall i\in\bar{\mathcal{I}}$ within the long tail of the non-selected indices is close to uniform and hence we can approximate its expectation by an average.
@@
Appendix E Evaluation statistics
In this section, we present again the experimental results from  Section 4.1; however, to demonstrate statistical significance, we show the error bars. This is important since every data point is aggregated using averaging across layers, heads, and test examples, as we will describe below. Therefore standard deviation of such an averaged metric is of interest.
[…]
Figures 16 and 17 show the evaluation on Human-eval dataset.  Figure 16 is similar to the Q&A plots as it shows how the reduction in the number of attention elements during prefill phase impacts the accuracy score (pass@1). Figure 17 focuses on the generative decoding phase and shows how the number of the needed V rows is affected by the elements that were selected in every row independently, which introduces the effect of GQA - since the indices of the selected elements are united across the heads in the group. The average and the standard deviation in this plot are taken across the following population of samples: number of test tasks $\times$ number of autoregressive forward passes $\times$ number of model layers $\times$ number of attention heads $\times$.

### v2

@@ 5 Related Work
A seminal approach in content-based sparsity is Top-k attention Gupta et al. (2021), which selects the $k$ largest elements per row of the attention score matrix. However, implementing efficient top-k selection is challenging in tiled kernel settings due to row-wise dependencies. Energon Zhou et al. (2022) addresses memory bandwidth by iteratively quantizing and filtering $\bm{K}$, but this restricts accurate computation of the final attention output, mirroring limitations seen when compensation is omitted in Top-{greek}j. Other methods, such as SpAtten Wang et al. (2021) and A3 Ham et al. (2020), require specialized hardware to accelerate top-k search, which Top-{greek}j avoids. SparQ Ribar et al. (2023) also employs top-k selection and introduces VMC-style compensation for omitted $\bm{V}$ rows, while the Swin Transformer Liu et al. (2021) adopts a fixed local sparsity pattern, restricting each token’s attention to its neighboring $M\times M$ patches.
Learned Token Pruning (LTP) Kim et al. (2022) is closely related to our work, leveraging a model-trained threshold to prune tokens in the attention matrix and achieving strong efficiency gains, but requiring full model retraining. In contrast, our approach identifies an inductive threshold within a pretrained model, enabling immediate deployment on large models without costly retraining or fine-tuning. Similarly,  Lou et al. (2024) select top-k elements but also depend on retraining. Methods like Sparsemax Martins and Astudillo (2016) and Entmax Peters et al. (2019) achieve sparsity via dynamic or iterative thresholding, but incur higher computational overhead due to sequence-wide sorting or iterative refinement. ReLA Zhang et al. (2021) induces sparsity through ReLU and normalization, but lacks sparsity control and is more expensive than static thresholding.
@@
Numerous works focused on fixed attention sparsity. For example, in Longformer Beltagy et al. (2020), only a few manually picked “global attention” tokens are allowed to attend to other tokens, whereas other tokens are restricted to the neighboring tokens in a dilated sliding window. Our work, on the other hand, capitalizes on identifying content-based sparsity in the attention matrix, as it offers higher accuracy gains. The work of Kim and Cho Kim and Cho (2021) requires retraining, although it accommodates dynamic shortening of the sequence on demand at inference time.
6 Limitations
This work targets attention sparsification and reduction of value-matrix accesses with a focus on maintaining accuracy. We do not assess wall-clock runtime improvements, as meaningful speedups require specialized, hardware-aware kernel optimizations, which are beyond this paper’s scope. Performance depends heavily on target hardware due to varying efficiency in handling sparse memory accesses, and we strongly believe this aspect warrants a separate in-depth study. Our experiments are limited to decoder-only transformer models within the LLaMA family, which cover common attention mechanisms such as GQA and MHSA that span across some of the most successful models (Mistral, QWEN, Phi, etc.). Extending the approach to larger or different transformer architectures and other attention methods remains an important direction for future work. Additionally, factors such as model size, sequence length, and attention patterns may influence the trade-offs between sparsity and accuracy in other settings.
@@ 7 Conclusion
We have presented Top-{greek}j, a new sparse attention algorithm based on fixed and calibrated thresholds. At inference time, it boils down to a simple elementwise operation, which overcomes the limitation of full-row dependent algorithms, such as Top-k, and thereby unconstrains tiling and distributed inference implementations. Moreover, only a small subset of calibration samples is required to calibrate the thresholds, which we showed to be resilient to distribution shifts. Therefore, a short calibration is needed once per model. Furthermore, we show minor to no performance degradation using 10× less attention elements at the computationally-bound prefill phase and using 3× (for shorter sequence tasks) to 10× less $\bm{V}$ matrix rows (for longer sequence tasks) at the memory bandwidth-bound generative decoding phase. These reductions unlock promising speedups for inference.
@@ Appendix A Impact Statement
This paper presents work whose goal is to advance the field of efficient Machine Learning. All potential societal consequences are mostly unrelated to the specific work but are more related to machine learning applications in general. One potential positive consequence of our work is that LLM technologies can be adopted by vendors and systems with lower resources, since our proposal unlocks deployment of LLMs in systems with lower memory bandwidth. We clearly and transparently state that our method is a lossy method (i.e., the accuracy of the downstream LLM task may degrade in favor of higher performance). Lossy deep learning models are a known practice in literature (e.g., when applying quantization and pruning), hence one should employ sufficient guardrails when deploying a lossy method in security-sensitive scenarios.
Appendix B Threshold Calibration Set Size
We have experimented with various calibration set sizes, and the main finding was that even with as few as 8 calibration samples, the model retains good accuracy. As the calibration set size grows larger, to a few hundred, the number of attention elements selected via thresholding ($\tilde{k}$) is approaching the desired $k$ on average, with a smaller variance. See Figure˜12 where we plot on the left the ratio between the effective number of elements that passed the threshold ($\tilde{k}$) and the desired $k$ (the closer to $1.0$ the better).
Figure 12: Calibration set size impact - LLaMA2-7B, calibration and evaluation on ARC-C. Left: the Y axis shows the average ratio $\tilde{k}/k$ (the closer to 1 the better the approximation of topk) and the X-axis (different bars) refers to different calibrations that were performed with calibration set sizes ranging from 8 to 380. Right: accuracy of LLaMA2-7b with Top-{greek}j on ARC-C dataset as a function of the calibration set size used to calibrate the thresholds.
Appendix C Threshold Calibration
In this appendix section, we show more calibrated threshold values (as a function of the sequence length) in more layers and heads.  Figure˜13 visualizes the thresholds as a function of a sequence length that was calibrated for The LLaMA2-7b model for the Hellaswag dataset. Different heads have different threshold values, hence the importance of per-head individual calibrations
@@ Appendix C Threshold Calibration
Appendix D Multi-k Cumulative Calibration
Algorithm˜1, which we presented in the paper, calibrates thresholds of a single attention head for a single target $k$ value. We propose to generalize it to calibrate all thresholds for a broader range of target $k$ and to do it in one pass over the calibration samples. We call such a calibration procedure “Multi-k Cumulative” (MKC).
@@ Appendix D Multi-k Cumulative Calibration
Algorithm˜2 describes how a threshold function $\theta(k)$ can be calibrated. First, for every given calibration sample, let $v$ represent the $r^{th}$ attention row. The $v$ undergoes a sorting and then is treated as a sequence of half-open intervals. For each interval, we treat its smaller endpoint as its threshold, and we also associate with it an effective $k$ (effective $k$ of a threshold w.r.t the vector $v$ is the number of elements that are greater than the threshold). For example, on line 2 we represent by $\langle r-1,[v_{1},v_{2})\rangle$ the interval between $v_{1}$ (inclusive) and $v_{2}$ (exclusive) that corresponds to the threshold $v_{1}$ and an effective $k$ of r1r-1. Second, per-calibration-sample interval sequences collected in the set $\Theta_{r}$ are merged to represent the function $\bar{k}(\theta):\mathbb{R}\rightarrow\mathbb{R}$ which maps each interval to a threshold and to an average effective $k$ achievable by this threshold on the calibration set. This merging is done by the MergeIntervals($\Theta_{r}$) subroutine, which computes the following:
@@
where we define $\Theta_{r,\theta}=\{t|\theta\in t_{1}\wedge t\in seq\wedge seq\in\Theta_{r}\}$ as a subset of intervals that include the $\theta$ in the interval. Note that the resulting average effective $k$ might be fractional due to the averaging. Finally, the desired $\theta(k)$ is given by the inverse of $\bar{k}(\theta)$, at points where the function $\bar{k}(\theta)$ maps to a natural value. The full implementation of the MergeIntervals routine is available in the supplied source code, and we show its visualization on Figure˜14.
@@
[…]
@@
3:  if is_prefill(XXthen
4:   $\bm{A}=Q(X)K^{T}(X)$
[…]
6:   for $r=1$ to n1n-1 do
[…]
@@
Figure˜15 illustrates the distribution of merged intervals ($\bar{k}(\theta)$) for selected attention layers and heads in the LLaMA2-7B model, following calibration via MKC on the Hellaswag dataset.
@@
Figure 15: MKC calibration of LLaMA2-7b on Hellaswag dataset. Right: the $\bar{k}_{r}(\theta)$ function as obtained from Algorithm˜2 by merging the intervals for sequence length $r=800$. Middle: zoom into a range of interest of $\bar{k}_{r}^{-1}(k)$ around $k=64$, the selected threshold is marked on using $\theta_{r}(64)=\bar{k}_{r}^{-1}(64)$, Left: thresholds obtained from MKC for all sequence lengths (also other than $r=800$.
It is worth mentioning that the main downside of the multi-k calibration is that it is very time and memory-consuming to perform all these interval mergings individually for every layer, head, and attention row id (i.e., sequence length). Speeding up the MergeIntervals routine, using reasonable approximations (such as subsampling) is an interesting research direction. The second slight downside is that it does not allow to apply topk at calibration, as we used to apply in Algorithm˜1, since during MKC there is a multitude of potential $k$ parameters we are targeting.
Appendix E V-Mean Compensation
In this section, we provide a formal proof for our $\bm{V}$-Mean Compensation (VMC) serving as a good approximation of the sparsified $\tilde{\bm{s}}\tilde{\bm{V}}$ product to the full $\mathbf{\bm{s}}\bm{V}$ product.
Lemma E.1.
Let $\bm{s}\in[0,1]^{n}$ denote the post-softmax attention vector, and let $\tilde{\bm{s}}\in[0,1]^{\tilde{k}}$ denote its thresholded variant, containing the $\tilde{k}$ selected values of $\bm{s}$, and let $\mathcal{I}\in\left\{0,\dots,n-1\right\}^{\tilde{k}}$ denote the indices of $\bm{s}$ that surpassed the threshold such that $\forall i\in\{0,\dots,\tilde{k}-1\}:\tilde{\bm{s}}_{i}=\bm{s}_{\mathcal{I}_{i}}$. Let $\bm{V}\in\mathbb{R}^{n\times d}$ be the value matrix and let $\tilde{\bm{V}}\in\mathbb{R}^{\tilde{k}\times d}$ be consisting only of the selected $\bm{V}$ rows such that $\forall i\in\{0,\dots,\tilde{k}-1\}:\tilde{\bm{V}}_{i}=\bm{V}_{\mathcal{I}_{i}}$. We claim that in the expectation, the full product $\bm{p}=\bm{s}\bm{V}\in\mathbb{R}^{d}$ is equal to the thresholded attention plus the residual probability mass $\beta$ (6) multiplied by the mean $\bm{V}$ row $\bm{\mu}$ (5). Namely,
$\forall 0\leq j<d:E[\bm{p}_{j}]=(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+\beta\bm{\mu}_{j}$ (9)
@@ Appendix E V-Mean Compensation
$\begin{split}E[\bm{p}_{j}]&=E\Big{[}\sum_{i=0}^{n-1}s_{i}\bm{V}_{ij}\Big{]}\\ &=E\Big{[}\sum_{i\in\mathcal{I}}s_{i}\bm{V}_{ij}\Big{]}+E\Big{[}\sum_{i\in\bar{\mathcal{I}}}s_{i}V_{ij}\Big{]}\\ &=(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+\sum_{i\in\bar{\mathcal{I}}}E[s_{i}\bm{V}_{ij}]\\ &\underset{s\perp\!\!\!\!\perp\bm{V}_{j}}{=}(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+\sum_{i\in\bar{\mathcal{I}}}E[s_{i}]E[\bm{V}_{ij}]\\ &\underset{s,\bm{V}_{j}\text{uniform}}{=}(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+\sum_{i\in\bar{\mathcal{I}}}\frac{\beta}{n-\tilde{k}}\bm{\mu}_{j}\\ &=(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+(n-\tilde{k})\frac{\beta}{n-\tilde{k}}\bm{\mu}_{j}\\ &=(\bm{\tilde{s}}\tilde{\bm{V}})_{j}+\beta\bm{\mu}_{j}\end{split}$ (10)
@@
    $s\perp\!\!\!\!\perp\bm{V}_{j}$, that is the attention vector $\bm{s}$ is statistically independent on the elements in the columns of matrix $\bm{V}$. They are conditionally independent given the input $X$ from which they were originally computed via $\bm{V}=\bm{XW_{V}}$.
@@
    The distribution of $s_{i},\forall i\in\bar{\mathcal{I}}$ within the long tail of the non-selected indices is close to uniform, and hence we can approximate its expectation by an average.
@@
Appendix F Evaluation statistics
In this section, we present again the experimental results from  Section˜4.1; however, to demonstrate statistical significance, we show the error bars. This is important since every data point is aggregated using averaging across layers, heads, and test examples, as we will describe below. Therefore standard deviation of such an averaged metric is of interest.
Figure˜16 focus on Q& tasks, showing the tradeoff between the model’s accuracy (y-axis) and the number of attention elements as selected by the Top-k or Top-{greek}j (x-axis). The standard deviation in the accuracy was provided by the LM-Eval evaluation harness, as a part of the standardized evaluation procedure. The average ratio presented on the x-axis and its standard deviation were computed over the following population: number of test samples $\times$ number of model layers $\times$ number of attention heads. To present the ratios, we first compute the absolute average and absolute standard deviation of the total count of attention elements and in the attention matrix, second - we normalize both the standard deviation and the average by the average number of attention elements when no sparsification took place.
Figures˜17 and 18 show the evaluation on Human-eval dataset.  Figure˜17 is similar to the Q&A plots as it shows how the reduction in the number of attention elements during prefill phase impacts the accuracy score (pass@1). Figure˜18 focuses on the generative decoding phase and shows how the number of the needed $\bm{V}$ rows is affected by the elements that were selected in every row independently, which introduces the effect of GQA - since the indices of the selected elements are united across the heads in the group. The average and the standard deviation in this plot are taken across the following population of samples: number of test tasks $\times$ number of autoregressive forward passes $\times$ number of model layers $\times$ number of attention heads $\times$.